OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = 3x + 2, what is f(g(x)) ?

OpenStudy (anonymous):

you put g in f's place like so

OpenStudy (anonymous):

f(g(x)) = 4(3x+2) +1

OpenStudy (anonymous):

12x+8 +1

OpenStudy (anonymous):

12x+ 9

OpenStudy (anonymous):

solve for zero: 12x = -9 x= -3/4

OpenStudy (anonymous):

i get it thanks so much!

OpenStudy (anonymous):

No problem :)

OpenStudy (accessdenied):

i dont quite understand why you solve for the zero if it only asked for f(g(x))

OpenStudy (accessdenied):

it seems sufficient to just say f(g(x)) = 12x + 9

OpenStudy (anonymous):

yepp, thats the answer. 12x+9

OpenStudy (anonymous):

I was just being overly cautious ;)

OpenStudy (anonymous):

do you guys know how to do this one? Let f(x) = 3x+1 and g(x) = x2 Find f(g(x)) then Find g(f(x)) and after Use complete sentences to explain any difference in part 1 and part 2.

OpenStudy (accessdenied):

x2 as in, x squared? it'd be the same process as the previous one i use x^2 to indicate the exponent so theres no confusion f(g(x)) = f(x^2) = 3(x^2) + 1 g(f(x)) = g(3x + 1) = (3x + 1)^2 im not very good with the worded explanation for this. i guess you'd write that the process is the same, the values you are applying it to varied

OpenStudy (anonymous):

ah alright thank you!